OpenStudy (anonymous):

write the equation of the line that is parallel to y=1/3x-3 and passes through point (6,2)

OpenStudy (anonymous):

The equation is already in the form of \(y=mx+c\), where \(m\) is the slope, so we know the slope is \(\frac{1}{3}\). Parallel lines have the same slope, so this is the slope of the new line as well. The formula for the equation of the line is \[y-y_{1} = m(x-x_{1})\] \(y-2 = \frac{1}{3}(x-6)\) using the point (6,2) \(3y-6=x-6\) \(3y=x\)

OpenStudy (anonymous):

so you just plug in the 6 and the 2?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

oooh okay thanks